OpenStudy (anonymous):

solving by addition: find X and Y 2x+9y=5 5x=6-3y

jhonyy9 (jhonyy9):

2x+9y=5 multiply by -5 and will get -10x-45y=-25 addition 5x+3y=6 multiply by 2 10x+6y=12 ---------------- 0 -39y=-13 y= -13/-39 y= 1/3 2x+9(1/3)=5 2x+9/3=5 2x+3=5 2x=2 x=1 x=1 y=1/3